OpenStudy (anonymous):

factor the quadratic expression completely, and find the roots of the expression. 1. 66x^2 - 130x + 56 2. 26x^2 - 159x - 65

OpenStudy (anonymous):

lol

myininaya (myininaya):

\[2\cdot 33x^2-2 \cdot 65x+2 \cdot 28\] each term has a two in common

OpenStudy (anonymous):

I've been working on these problems for 24 hrs.

myininaya (myininaya):

\[2(33x^2-65x+28)\]

OpenStudy (mathteacher1729):

Supaman -- what are you finding difficult about these problems? Where do you get stuck?

OpenStudy (anonymous):

the answer i got is 2(11x7)(3x-4)

OpenStudy (anonymous):

now wonder. i would cheat of course that will not help on a test. myininaya has a nice method. maybe will share http://www.wolframalpha.com/input/?i=66x%5E2+-+130x+%2B+56+

OpenStudy (anonymous):

im trying to see if my answer is right. if so how the heck do i do the root

OpenStudy (anonymous):

that's the answer: 2(11x-7)(3x-4)

OpenStudy (anonymous):

that web site i so dont understand.

myininaya (myininaya):

now see if you factor \[33x^2-65x+28\] by finding two factors of 33(28) that have product 33(28) and sum -65 so 33(28)=66(14)=22(42)=44(21)=-44(-21)

myininaya (myininaya):

-65=-44-21

myininaya (myininaya):

-65x=-44x-21x

myininaya (myininaya):

33x^2-44x-21x+28

OpenStudy (anonymous):

um? im lost sir. like i said i got 2(11x7(3x-4) is that right or not?

OpenStudy (anonymous):

yes

myininaya (myininaya):

\[11x(3x-4)-7(3x-4)=(3x-4)(11x-7)\]

OpenStudy (anonymous):

your answer is right

myininaya (myininaya):

but don't forget about the 2 2(3x-4)(11x-7)

OpenStudy (anonymous):

JESUS! so how do i finish? it thats where im left off at.

OpenStudy (anonymous):

that is it, you are done

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

you just have to find the roots

OpenStudy (anonymous):

so thats the root too?

OpenStudy (anonymous):

do you know the definition of a root?

OpenStudy (anonymous):

yea thats what i was wondering how to do.

myininaya (myininaya):

find the roots by finding where the expression is 0

OpenStudy (anonymous):

graphically, a root is where the function crosses the x axis.

OpenStudy (mathteacher1729):

Quick question Supaman -- can you expand these binomials? (x-3)(x+4) = (2x + 3)(4x-1) = (4x -2)(7x+3) =

OpenStudy (anonymous):

nope i hate3 math im failing and im on my last week in college.. math is a subject i can never use or do.

OpenStudy (anonymous):

i dont know what that means mathteacher.

myininaya (myininaya):

hes asking if you know how to multiply binomials

OpenStudy (mathteacher1729):

>i dont know what that means mathteacher. Ok, that makes sense then. You have to go and learn how to multiply binomials before you can learn to factor them. As for math not being useful in life... managing your own money is all about math. (not necessarily factoring polynomials though...) :-p

OpenStudy (mathteacher1729):

Here is Paul's Online Math Notes, which feature very nicely written explanations of all sorts of math topics. This is their Factoring Polynomials lesson: http://tutorial.math.lamar.edu/Classes/Alg/Factoring.aspx

OpenStudy (anonymous):

right. so, how do i finish this??

myininaya (myininaya):

didn't you already finish the first one?

myininaya (myininaya):

2(3x-4)(11x-7)

OpenStudy (anonymous):

yes, but i thought i have to root, and @teacher can't open that link it says some off stuff.

myininaya (myininaya):

yes just find where the expression is 0 to find the roots

OpenStudy (mathteacher1729):

The problems you posed here are advanced polynomial factoring. First, you have to learn how to add, subtract and multiply polynomials: http://tutorial.math.lamar.edu/Classes/Alg/Polynomials.aspx Then after that, learn how to factor them: http://tutorial.math.lamar.edu/Classes/Alg/Factoring.aspx After those two lessons, you should be ready to try the problems you posted here.

myininaya (myininaya):

2(0)(11x-7)=0 so when 3x-4=0 we have a root 2(3x-4)(0)=0 so when 11x-7=0 we have a root

OpenStudy (anonymous):

do i 2(11(0))(3(0)-4)